OpenStudy (anonymous):

vector (2,-4) & (1,-3,2) find the lengths

OpenStudy (anonymous):

ok. distance formula is: square each component. sum the results. take square root. It's the same whether vectors are 2D, 3D, etc..

OpenStudy (anonymous):

So for the first vector: \[\sqrt{2^2+(-4)^2}=\sqrt{20}=2\sqrt5\]

OpenStudy (anonymous):

thank you a thousand times

OpenStudy (anonymous):

no prob. What did you get for the second vector?

OpenStudy (anonymous):

sqr root 14

OpenStudy (anonymous):

if thats not right i've made a mistake factoring

OpenStudy (anonymous):

looks correct.